OpenStudy (anonymous):

The mean of a set of x numbers is 17. When 3x is added to that data set, the new mean is 18. Find x.

jimthompson5910 (jim_thompson5910):

S/x = 17 ----> S = 17x (S+3x)/(x+1) = 18 (17x+3x)/(x+1) = 18 (20x)/(x+1) = 18 20x = 18(x+1) 20x = 18x+18 20x - 18x = 18 2x = 18 x = 18/2 x = 9

OpenStudy (anonymous):

Thanks jim. Makes sense.

jimthompson5910 (jim_thompson5910):

You're welcome

OpenStudy (anonymous):

since mean of x numbers is 17 => Sum of x numbers = 17x we add one more number i.e. 3x so now the numbers are (x + 1) and sum of (x + 1) = 17x + 3x = 20x 20x so mean of (x + 1) numbers = ------ x + 1 ATQ, 20x ----- = 18 => 20x = 18(x + 1) => 20x = 18x + 18 => 20x - 18x = 18 x + 1 => 2x = 18 => x = 18/2 => x = 9